A rectangular parking lot is enclosed by 400 feet of fencing. The longer sides are three times as long as the shorter sides. What is the area of the lot?

Answer by rothauserc(4718)

let l be length and w be width, then
l = 3w
Perimeter(P) = 2l + 2w
400 = 2(3w) + 2w
400 = 8w
w = 50
l = 3*50 = 150
Area = 50 * 150 = 7500 square feet